OpenStudy (anonymous):

2=-1/3(4)+b, how do i get the answer?

OpenStudy (anonymous):

2 = -1/3(4) +b or is b part of the denominator

OpenStudy (anonymous):

b = 3.33

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2 = [-1/3 * 4 ] +b 2 = -4/3 +b 2 + 4/3 =b 10/3 = b which is equal to 3.33 recurring

OpenStudy (anonymous):

Than you so very much. That is the answer and I understand how you got that. Thank you :0)

OpenStudy (anonymous):

your welcome